OpenStudy (anonymous):

33. What is the slope of the line that passes through the pair of points?(–2, 7), (18, 1)

OpenStudy (anonymous):

You use the slope formula: m = (y2 - y1)/(x2 - x1) m = (7 - 1) / (-2 - 18) = -6/20 = -3/10